From a pack of card two cards are drawn
find the probability that 1. both are black
1
Expert's answer
2020-10-12T18:15:08-0400

Condition probability. A deck of cards has 26 black card.

"P(B_1)=\frac{26}{52}=\frac{1}{2}"

"P(B_2/B_1)=\frac{25}{51}"

"P(B_1B_2)=\frac{1}{2}×\frac{25}{51}=\frac{25}{102}"